# KID-PPG: Knowledge Informed Deep Learning for Extracting Heart Rate from a Smartwatch

### v1

@@ 2 Methodology
These are incorporated into our DL models through three mechanisms: Explicit Source Separation, Guided Probabilistic Inference and Data Augmentation. An overview of our methodology is presented in Fig. 1.
@@ 2.1 Motion Artifact Removal
As is common practice, HR is inferred on 8-second windows with a 2-second overlap [14, 13, 12, 16, 15, 19, 4, 5, 20, 21]. Hence, the i-th 8-second PPG sample is denoted as the $(N\times 1)$ vector $\textbf{x}_{ppg_{i}}=[x_{ppg}(t_{i}),x_{ppg}(t_{i}+\Delta t),...,x_{ppg}(t_{i}+8)]$, where $\Delta t=\frac{1}{f_{s}}$, $f_{s}$ is the sensor’s sampling frequency, and $N=8sec\cdot f_{s}$ samples. Similarly, the corresponding BVP is denoted as $\textbf{x}_{bvp_{i}}$. The 3-axis ACC samples form the $(N\times 3)$ matrix $X_{acc_{i}}=[\textbf{x}_{acc_{x_{i}}},\textbf{x}_{acc_{y_{i}}},\textbf{x}_{acc_{z_{i}}}]$. Robust HR inference depends on the heartbeat-related $\textbf{x}_{bvp}$ [4]. However, MA can significantly distort the morphology of BVP in the PPG observation (Eq. 1) [5]. Many SP methods employ source separation to approximate the unmixing process $f_{mix}^{-1}$, allowing the HR inference module to rely on an approximation $\hat{x}_{bvp}(t)\approx x_{bvp}(t)$ [5, 4, 19, 3], thus mitigating the effect of MA.
@@ 2.3 Guided Probabilistic Heart Rate Extraction
where $F_{\mu_{hr},\sigma_{hr}}(x)$ is the cumulative distribution of the Gaussian distribution $\mathcal{N}(\mu_{hr},\sigma_{hr}^{2})$: $F_{\mu_{hr},\sigma_{hr}}(x)=\Phi(\frac{x-\mu_{hr}}{\sigma_{hr}})$. This error classifier predicts a high probability of the HR estimation being untrustworthy (error $\geq Thr$) when $p_{error}\leq 0.5$.
@@
Figure 4: Inference after filtering the BVP component out of the PPG. Example taken from S6 of PPGDalia. (a) Inferences across the entire session from Q-PPG (red), probabilistic (blue) and guided probabilistic (orange). For the probabilistic models, the range of one standard deviation is also presented. (b) Sample example with initially clean PPG (grey) and synthetically degraded (black). (c) Frequency domain representation of the example sample and HR inferences from the three models. True HR is presented with a green circle.
@@ 2.4 Data Augmentation
We propose a data augmentation scheme to address the limited number of high HR samples in the available datasets, similarly to [17]. We generate synthetic PPG waveforms corresponding to higher heart rate frequencies, forming the high-heart-rate training subset. The following procedure is followed:
@@ 3 Experimental Setup
For our experiments we use the publicly available PPGDalia data [12]. This dataset comprises synchronized ECG and wristworn PPG and acceleration recordings from 15 subjects, with approximately two-hour recording sessions per subject. A systematic temporal shift between PPG and ACC was identified and manually corrected. The ECG-based HR provided in the dataset serves as the ground truth. We adopt the leave-one-subject-out cross-validation procedure proposed in [12] for all experiments.
@@
The probabilistic models where trained using the negative log-likelihood (NLL) loss and the same training strategy used for the non-probabilistic ones. NLL is also used to evaluate their performance. Furthermore, we use the True Positive Rate (TPR) and the F1 score to evaluate the classifier $p_{error}$ for its ability to correctly identify untrustworthy samples. We consider a positive classification when the error classifier predicts a high probability of error, $p_{error}(Thr)\leq 0.5$.
@@ 4.1 Results on Motion Artifact Removal
The adaptive linear filter effectively removes MA that are linearly coupled with acceleration signals. An illustrative example is provided in Fig. 6 for subjects S6 during an episode where the subject is walking up and down stairs. The DL model alone fails to decouple motion information, resulting in significant errors (activity MAE 13.66 BPM). In contrast, the Adaptive + DL Model combination substantially reduces the error (activity MAE 2.74 BPM). In particular, during periods when Q-PPG loses track of the HR, the BVP signal, located at $hr$ and $2\cdot hr$, remains discernible. The loss of HR tracking is not attributed to severe degradation of BVP due to MA.
@@
Figure 7: Effect of the MA Removal for the stairs activity of subjects S5 (top), and S9 (bottom) with Q-PPG (blue) and Adaptive + Q-PPG (orange). Ground truth HR is represented in green.
@@
Adaptive + Q-PPG 3.80 3.50 2.07 5.18 5.76 3.38 1.59 6.99 8.94 2.75 3.37 5.61 1.49 2.45 2.80 3.98
Adaptive + Attention 4.55 3.12 1.88 5.27 4.00 2.54 1.28 7.20 8.60 2.61 3.06 4.91 1.50 2.28 2.64 3.70
Adaptive + Attention + HR Aug 4.27 3.46 2.07 5.61 3.01 2.74 1.39 7.13 9.53 2.77 3.58 4.52 1.48 2.48 2.84 3.79
@@
KID-PPG ($Thr=10BPM$) 2.99 2.54 1.84 3.83 2.8 2.56 1.50 5.16 4.67 2.33 2.90 3.73 1.72 2.33 1.91 2.85
Table 2: Mean Absolute Error performance of point-estimate HR and probabilistic (after retention) models on the PPGDalia dataset. All values are in Beats-per-Minute.
@@ 4.3 Guided Probabilistic Inference
The NLL losses for the probabilistic models are summarized in Table 3. Our proposed methodology, KID-PPG, significantly outperforms BeliefPPG [15] (NLL 4.78) and reduces NLL by approximately 40%.
@@
Figure 9: Example of probabilistic output of augmented (orange) vs non-augmented model (blue) for subject S9 and ground truth HR (green circle). For the example, its frequency domain representation verifies that the PPG signal is considerably corrupted.
@@ 5 Discussion
Catastrophic MA remains a significant challenge, particularly evident in subjects S8 and S9, which consistently exhibit high MAE. Despite various proposed solutions, no one has achieved a low MAE for these two subjects. Furthermore, the rejection of a significant portion of their samples by KID-PPG (27.48% and 53.16% of their samples respectively $\left(Thr=10BPM\right)$) indicates severe BVP corruption. This underscores the need for a probabilistic HR estimator and a robust error classifier based on the morphology of the BVP components.
@@
In integrating BVP-morphology-related prior knowledge into KID-PPG, we have proposed a method for creating realistic PPG signals with extreme BVP degradation. Exploring alternative approaches to generate synthetic MA-affected PPG samples and expanding the understanding of MA dynamics with the BVP could inform the design of more realistic augmentation techniques. Evaluating MA degradation presents an additional challenge, as HR-inference accuracy may not fully reflect the level of degradation. Addressing these challenges will be crucial to advancing the robustness and reliability of PPG-based HR inference systems.

### v2

@@ 2 Methodology
This prior-knowledge is incorporated into our DL models through three mechanisms: Explicit Source Separation, Guided Probabilistic Inference and Data Augmentation. An overview of our methodology is presented in Fig. 1.
@@ 2.1 Motion Artifact Removal
As is common practice, HR is inferred on 8-second windows with a 2-second overlap [14, 13, 12, 16, 15, 19, 4, 5, 20, 21]. Hence, the i-th 8-second PPG sample is denoted as the $(N\times 1)$ vector $\textbf{x}_{ppg_{i}}=[x_{ppg}(t_{i}),x_{ppg}(t_{i}+\Delta t),...,x_{ppg}(t_{i}+8)]$, where $\Delta t=\frac{1}{f_{s}}$, $f_{s}$ is the sensor’s sampling frequency, and $N=8sec\cdot f_{s}$ samples. Similarly, the corresponding BVP is denoted as $\textbf{x}_{bvp_{i}}$. The 3-axis acceleration (ACC) samples form the $(N\times 3)$ matrix $X_{acc_{i}}=[\textbf{x}_{acc_{x_{i}}},\textbf{x}_{acc_{y_{i}}},\textbf{x}_{acc_{z_{i}}}]$. Robust HR inference depends on the heartbeat-related $\textbf{x}_{bvp}$ [4]. However, MA can significantly distort the morphology of BVP in the PPG observation (Eq. 1) [5]. Many SP methods employ source separation to approximate the unmixing process $f_{mix}^{-1}$, allowing the HR inference module to rely on an approximation $\hat{x}_{bvp}(t)\approx x_{bvp}(t)$ [5, 4, 19, 3], thus mitigating the effect of MA.
@@ 2.3 Guided Probabilistic Heart Rate Extraction
where $F_{\mu_{hr},\sigma_{hr}}(x)$ is the cumulative distribution of the Gaussian distribution $\mathcal{N}(\mu_{hr},\sigma_{hr}^{2})$: $F_{\mu_{hr},\sigma_{hr}}(x)=\Phi(\frac{x-\mu_{hr}}{\sigma_{hr}})$ and $Thr$ is the threshold above which the error is considered significant. This error classifier predicts a high probability of the HR estimation being untrustworthy (error $\geq Thr$) when $p_{error}\geq CL$, where $CL$ is the required confidence level. As is usual in classifiers we set the confidence level to 0.5, although it can be freely selected to stricter values depending on the application.
@@
To improve the robustness of our error classifier, we guide the network to base its HR inference on the BVP component, as relying on HR-inference loss only can lead to spurious behavior. Fig. 4 illustrates this with a synthetic example generated from PPGDalia (S6). Here we have manually removed the BVP component, by applying a bandstop filter with cut-off frequencies around $HR$, $2\cdot HR$ and $3\cdot HR$, yet both the probabilistic and the Q-PPG models continue to infer HR (MAE 4.11 BPM and 3.33 BPM accordingly). The probabilistic model is overconfident, with the $p_{error}(Thr=10BPM)$ classifier dropping only $7.44\%$ of the samples.
@@
Figure 4: Inference after filtering the BVP component out of the PPG. Example taken from S6 of PPGDalia. (a) Inferences across the entire session from Q-PPG (red), probabilistic (blue) and guided probabilistic (orange). For the probabilistic models, the range of one standard deviation is also presented. (b) Sample example with initially clean PPG (grey) and synthetically degraded (black). (c) Frequency domain representation of the example sample and HR inferences from the three models. True HR is presented with a green circle. Both Q-PPG and probabilistic models estimate HR close to the ground truth, indicating potential learned shortcuts since there is no heart rate information in the signals. In contrast, the guided probabilistic model estimates a large standard deviation identifying the lack of relevant information.
@@ 2.4 Data Augmentation
We propose a data augmentation scheme to address the limited number of high HR samples in the available datasets, similarly to [17]. After removing the MA with $f_{mix}$, from the samples of the training set, we generate synthetic PPG waveforms corresponding to higher heart rate frequencies, forming the high-heart-rate training subset. The following procedure is followed:
@@ 3 Experimental Setup
For our experiments we use the publicly available PPGDalia data [12]. This dataset comprises synchronized ECG and wrist-worn PPG and acceleration recordings from 15 subjects, with approximately two-hour recording sessions per subject. The PPG and acceleration signals were collected using the Empatica E4 wristband. During the two-hour session, the subjects underwent diverse activities to simulate daily life conditions: resting, ascending/descending stairs, playing table soccer, cycling, driving a car, having lunch, walking and working in an office. Between the activities there is a transition period. A systematic temporal shift between PPG and ACC was identified and manually corrected. The ECG-based HR provided in the dataset serves as the ground truth. We adopt the leave-one-subject-out cross-validation procedure proposed in [12] for all experiments.
Additionally, to showcase KID-PPG’s generalizability we validate a pre-trained model on the WESAD dataset [32]. WESAD is comprised of recordings from 15 subjects. The session for each subject involves activities evoking various stress levels. More details on the activities can be found in the original publication [32]. We select a model trained during the PPGDalia cross-validation process and validate it on all the subjects of WESAD without any further training.
@@
The probabilistic models where trained using the negative log-likelihood (NLL) loss and the same training strategy used for the non-probabilistic ones. NLL is also used to evaluate their performance. Furthermore, we use the True Positive Rate (TPR) and the F1 score to evaluate the classifier $p_{error}$ for its ability to correctly identify untrustworthy samples. We consider a positive classification when the error classifier predicts a high probability of error, $p_{error}(Thr)\leq 0.5$, the error threshold was arbitrarily selected at $Thr=10BPM$.
@@
WESAD
• KID-PPG A pretrained model from the PPGDalia experiments is validated on WESAD without any further training.
@@ 4.1 Results on Motion Artifact Removal
The adaptive linear filter effectively removes MA that are linearly coupled with acceleration signals. An illustrative example is provided in Fig. 6 for subjects S6 during an episode where the subject is walking up and down stairs. The DL model alone fails to decouple motion information, resulting in significant errors (activity MAE 13.66 BPM). In contrast, the Adaptive + DL Model combination substantially reduces the error (activity MAE 2.74 BPM). In particular, during periods when Q-PPG loses track of the HR, the BVP signal, located at $hr$ and $2\cdot hr$, remains discernible and thus the loss of HR tracking is not attributed to severe degradation of BVP due to MA.
@@
Figure 7: Effect of the MA Removal for the stairs activity. Top: Subjects S5 with Q-PPG (blue) and Adaptive + Q-PPG (orange). The ground truth heart rate is clearly visible in the spectrogram as a black line as indicated by the yellow arrow. Bottom: Subject S9. Ground truth HR is represented in green.
@@
Adaptive + Q-PPG 3.80 (6.77) 3.50 (6.99) 2.07 (4.07) 5.18 (7.67) 5.76 (14.92) 3.38 (7.24) 1.59 (2.15) 6.99 (9.79) 8.94 (10.86) 2.75 (5.80) 3.37 (7.40) 5.61 (8.85) 1.49 (2.72) 2.45 (4.68) 2.80 (5.87) 3.98 (7.05)
Adaptive + Attention 4.55 (9.22) 3.12 (5.97) 1.88 (3.54) 5.27 (7.73) 4.00 (12.14) 2.54 (5.83) 1.28 (1.85) 7.20 (10.41) 8.60 (11.07) 2.61 (6.03) 3.06 (6.39) 4.91 (9.10) 1.50 (2.59) 2.28 (4.99) 2.64 (6.31) 3.70 (6.88)
Adaptive + Attention + HR Aug 4.27 (7.68) 3.46 (8.83) 2.07 (3.78) 5.61 (9.78) 3.01 (9.41) 2.74 (6.99) 1.39 (2.31) 7.13 (9.65) 9.53 (14.50) 2.77 (6.36) 3.58 (7.76) 4.52 (7.16) 1.48 (2.75) 2.48 (5.32) 2.84 (5.82) 3.79 (7.21)
@@
KID-PPG ($Thr=10BPM$) 2.99 (4.21) 2.54 (4.40) 1.84 (2.59) 3.83 (5.48) 2.8 (6.52) 2.56 (5.33) 1.50 (1.84) 5.16 (7.03) 4.67 (6.92) 2.33 (3.77) 2.90 (4.80) 3.73 (5.13) 1.72 (2.28) 2.33 (3.61) 1.91 (2.79) 2.85 (4.45)
KID-PPG (retention) 79.89 87.08 92.73 66.76 88.35 93.12 99.33 72.51 48.32 86.63 86.12 79.25 99.16 92.57 83.55 83.69
Table 2: Mean Absolute Error and Standard Deviation of the Absolute Error (in parenthesis)
performance of point-estimate HR and probabilistic (after retention) models on the PPGDalia dataset. All values are in Beats-per-Minute with the exception of the retention percentages which are presented for KID-PPG (probabilistic model).
@@ 4.3 Guided Probabilistic Inference
The NLL losses for the probabilistic models are summarized in Table 3. Our proposed methodology, KID-PPG, significantly outperforms BeliefPPG [15] (NLL 4.78) and reduces NLL by approximately 40%. KID-PPG also manages to generalize to a never-before-seen dataset (WESAD) achieving an NLL of 3.15, outperforming BeliefPPG (NLL 4.7 - 32% reduction).
@@
Figure 9: Example of probabilistic output of augmented (orange) vs non-augmented model (blue) for subject S9 and ground truth HR (green circle). For the example, its frequency domain representation verifies that the PPG signal is considerably corrupted. The augmented probabilistic model manages to identify the lack of BVP content, presenting a high standard deviation, in contrast to the non-augmented one.
@@ 5 Discussion
Catastrophic MA remains a significant challenge, particularly evident in subjects S8 and S9, which consistently exhibit high MAE. Despite various proposed solutions, no one has achieved a low MAE for these two subjects (Table 2). Furthermore, the rejection of a significant portion of their samples by KID-PPG (27.48% and 53.16% of their samples respectively $\left(Thr=10BPM\right)$) indicates severe BVP corruption. This underscores the need for a probabilistic HR estimator and a robust error classifier based on the morphology of the BVP components.
@@
In integrating BVP-morphology-related prior knowledge into KID-PPG, we have proposed a method for creating realistic PPG signals with extreme BVP degradation. Exploring alternative approaches to generate synthetic MA-affected PPG samples and expanding the understanding of MA dynamics with the BVP could inform the design of more realistic augmentation techniques. Evaluating MA degradation presents an additional challenge, as HR-inference accuracy may not fully reflect the level of degradation, for example Fig.4c. Addressing these challenges will be crucial to advancing the robustness and reliability of PPG-based HR inference systems.